graphing:
use slopes to determine if points (-5,-2)(-3,1)(3,0)(5,3) are the vertices of a parallelogram

Answer by venugopalramana(3286)

use slopes to determine if points A=(-5,-2)...B=(-3,1)...C=(3,0)...D=(5,3) are the vertices of a parallelogram
SLOPE OF AB=(1+2)/(-3+5)=3/2
SLOPE OF CD=(3-0)/(5-3)=3/2....HENCE AB||CD
SLOPE OF AC=(0+2)/(3+5)=2/8=1/4
SLOPE OF BD=((3-1)/(5+3)=2/8=1/4....HENCE AC||BD
HENCE ABCD IS PARALLELOGRAM